The temperature coefficient for a reaction is 2. The increase in rate as a result of temperature rise from 10°c to 100°c is

Grade:11

2 Answers

Arun
6 years ago
Dear Abhinav
 
\frac{k_T_2}{k_T_1}=\mu ^\frac{{T_2-T_1}}{10} =\mu ^{\frac{\Delta T}{10}}
 
 = 2^(90/10)
 = 2^9
now you can simply find the increase in rate of reaction
 
Regards
Arun(askIITians forum expert)
ankit singh
3 years ago
Temperature coefficient is defined as change in rate ratio by an increase in 10^{o}C.
Mathematically,     Temperature coefficient = \frac{k_{t+10^{o}}}{k_{t}}
Therefore, in the given situation we can conclude that when temperature changes from 30^{o}C to 40^{o}C then it is given that temperature coefficient is 2.
Hence, when temperature changes from 30^{o}C to 100^{o}C then rate of reaction will be 2^{7} which is equal to 128.